17/2 + 3½ = 17/2 + 3½ = 17/2 + 3½ = 17/2 + 3½ = 17/2 + 3½ =

 17/2 + 3½ =                   
 17/2 + 3½ =                    
 
17/2 + 3½ =                    
17/2 + 3½ =                    
17/2 + 3½ =                    
 
 
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

17/2 + 3 whole 1⁄2 = 12.
First we have to convert mixed fraction into improper fraction = 3 whole1⁄2
= 7/2
Fractions with same denominators are called like fractions.

Sum of two like fractions = (17/2) + (7/2) = (17 + 7)/2

= 24/2



Thanks